OpenStudy (anonymous):

if a=18 and b=3, find the value of a/b

OpenStudy (zzr0ck3r):

\[\frac{ 18 }{ 3 }\]?

OpenStudy (zzr0ck3r):

hint: 6 * 3 = 18

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

6--- bc 18/3 = 6

OpenStudy (anonymous):

just plug in the numbers